Mary has 30 hats. She has 5 green, 7 blue, 12 red, and 6 brown. If she selects two hats at random, what is the probability that both hats are different colors? Round your answer to two decimal places.
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

 
Hi
Mary has 30 hats. She has 5 green, 7 blue, 12 red, and 6 brown.
selects two hats at random
P(they are different colors) = 1 - P(both are the same color)
P(both are the same color) = 5/30*4/29 + 7/30*6/29 + 12/30*11/29 + 6/30*5*29
= (20 + 42 + 132 + 30)/30*29 = 224/30*29 = .2575
P(they are different colors) = 1 -.2575 = .7425

Answer by stanbon(75887)

Mary has 30 hats. She has 5 green, 7 blue, 12 red, and 6 brown. If she selects two hats at random, what is the probability that both hats are different colors? Round your answer to two decimal places.
---------------
Ans: 1 - P(hats are same color)
----
= 1 - [5C2 + 7C2 + 12C2 + 6C2]/30C2
-----
= 1 - [112/435]
-----
= 323/435
----
=0.743
==========
Cheers,
Stan H.
==========